OpenStudy (anonymous):

please! check my anwser. Four cards drawn from standard desk of cards .Each card is replaced before the next one is drawn .Find eanh probability

OpenStudy (chaise):

The probability of one type of card (red, ace of diamond) is 1/52.

OpenStudy (anonymous):

P ( 4 jacks) = (4/52) + (4/52) + (4/52) + (4/52) P( at most 1 jack ) = ?

OpenStudy (dumbcow):

Drawing a jack is an independent event, multiply probabilities of independent events P(4 jacks) = (4/52)(4/52)(4/52)(4/52) P(at most 1 jack) = P(0 jacks) + P(1 Jack) P(0 jacks) = (48/52)(48/52)(48/52)(48/52) P(1 jack) = (4C1)(4/52)(48/52)(48/52)(48/52)

OpenStudy (dumbcow):

its similar to binomial theorem n = 4 p = 4/52

OpenStudy (anonymous):

for jacks means jack and jack and jack and jack. since the cards are replaced it is \[(\frac{4}{52})^4\] oh, what dumbcow said, never mind

OpenStudy (anonymous):

can you help me P( exactly 3 jacks) ?

OpenStudy (anonymous):

(4/52)^4 ?

OpenStudy (anonymous):

yess exactly 3 jacks

OpenStudy (anonymous):

3 jacks, one not jack \[(\frac{4}{52})^3\times \frac{48}{52}\times 4\]

OpenStudy (anonymous):

4= ?

OpenStudy (anonymous):

four cards. jack, jack, jack, not jack jack, jack, not jack, jack jack, not jack, jack, jack not jack, jack, jack, jack

OpenStudy (anonymous):

that explains why you multiply by 4

OpenStudy (anonymous):

each one of these has the same probability, each one is \[(\frac{4}{52})^2(\frac{48}{52})\]

OpenStudy (anonymous):

but there are 4 disjoint possibilities, hence the multiply by 4 part